A cheerleader standing at the top of a human pyramid has 1801 J of gravitational potential energy. If the
cheerleader is 4.50 m above the surface of the earth, her mass must be
a) 20.4 kg
b) 31.7 kg
c) 40.8 kg
d) 95.4 kg
1
Expert's answer
2020-05-25T10:58:07-0400

The gravitational potential energy of cheerleader

"PE=mgh"

Hence, the mass of cheerleader

"m=\frac{PE}{gh}=\frac{1801\:\rm J}{9.8\:\rm N/kg\times 4.50\: m}=40.8\:\rm kg"

Answer: c)